OpenStudy (anonymous):

prove this identity 3cos2x=2-sinx

OpenStudy (xapproachesinfinity):

there is something wrong! the right hand side shouldn't be that way

OpenStudy (xapproachesinfinity):

hmm 3cos2x=3(1-2sin^2x)=3-6sin^2x

OpenStudy (xapproachesinfinity):

if you try half angle identity it won't give that result as well

OpenStudy (anonymous):

RHS does not equal the lLHS